OpenStudy (anonymous):

A(-1,1), B(3,3), C(6,2) lie on the circumference of a circle. Find the Center of the circle.

OpenStudy (anonymous):

The equation of a circle centered at (h, k) is given by \[(x-h)^2+(y-k)^2=r^2\] Given the points A, B, and C, you have the following system of equations: \[\begin{cases}(-1-h)^2+(1-k)^2=r^2\\ (3-h)^2+(3-k)^2=r^2\\ (6-h)^2+(2-k)^2=r^2\end{cases}\]Think you can solve that? (By the way, I haven't done a problem like this in a long time, so there may or may not be an easier way.)

OpenStudy (anonymous):

Thanks a lot for the help I'll get back with the answer.

OpenStudy (anonymous):

Ive looked out but it is apparently a straight line questions

OpenStudy (ash2326):

Which equation is straight line's ?